OpenStudy (anonymous):

I need help on this word expression problem. The product of 4 and a number is 18 more than twice the number. This can be simplified (I think) to 4n - 18 = 2n. While a guess or straight answer can be typed, I'd prefer an explanation on the steps. :)

OpenStudy (paxpolaris):

your equation is correct.

OpenStudy (anonymous):

Alright, can you help me solve n?

OpenStudy (paxpolaris):

alright let's move all the term with n on one side of the equation ...

OpenStudy (paxpolaris):



OpenStudy (paxpolaris):



OpenStudy (anonymous):

Okay.

OpenStudy (paxpolaris):

4n -2n -18 = 0

OpenStudy (paxpolaris):

so thats: 2n -18 = 0

OpenStudy (paxpolaris):

now lets move 18 to the right side of the equation -18 will become plus 18 2n = 0 +18

OpenStudy (paxpolaris):

2n =18 lastly you divide both sides by 2

OpenStudy (paxpolaris):

so, n = 9 gtg

OpenStudy (anonymous):

Thanks!